OpenStudy (anonymous):

discuss the steps that are necesarry for mult and dividing rational expressions

OpenStudy (anonymous):

and come up with an example

OpenStudy (dumbcow):

multiplying --> multiply numerators , then denominators by distributing dividing --> first flip the 2nd fraction, then multiply example: \[\frac{x}{x+3} \div \frac{x-1}{2x+3}\] flip 2nd fraction and multiply \[\frac{x}{x+3} * \frac{2x+3}{x-1}\] distribute both numerator and denominator \[\rightarrow \frac{2x^2 +3x}{x^2 +2x-3}\]

OpenStudy (anonymous):

THANK YOU!!

OpenStudy (dumbcow):

yw :)

OpenStudy (anonymous):

could you possibly do the same exact thing for add and subtracting @dumbcow

OpenStudy (anonymous):

@dumbcow (: